# How Not to Detect Prompt Injections with an LLM

### v2

@@ 5.4 RQ3: Fine-tuned vs Base FNR
Design. We quantitatively measure the advantage (if any) of the fine-tuning process used in Strong KAD defenses such as DataSentinel by comparing the false negative rate (FNR) of the detection model $g$ with its corresponding base model Mistral-7B, under two adversarial settings: Combined Attack and DataFlip.
@@
Setup. We consider contaminated inputs crafted using Combined Attack and DataFlip on the Prompt Injection dataset from Section 5.1.2. We report FPRs and FNRs (Section 5.1.3) corresponding to each attack, referred to as FNRCombined and FNRDataFlip, for the fine-tuned detection LLM and base Mistral.
Results. We see in Table 2 that, while fine-tuning on adversarial samples similar to the Combined Attack dramatically improves FPR, the FNRCombined of the DataSentinel detector is only marginally better than the base Mistral model. Furthermore, we find that FNRDataFlip of the finetuned detector is significantly worse than the base model, with differences as large as $80\%$ for most injected tasks.
@@
In fact, we note that the values of FNRCombined and FNRDataFlip both indicate a high tendency to follow the injected instruction as per our analysis in Section 3. The Combined Attack takes no advantage of the fact that the detection LLM is attempting to follow the instruction, resulting in the secret key not being produced. However, when we switch to DataFlip, instruction following results in the extraction of the secret key, yielding high FNR. This once again underscores the flawed nature of the KAD objective and the negative impact of Strong KAD defenses, which optimize for it.
6 Discussions
@@
In this section, we discuss the limitations of potential attempts to improve KAD (Section 6.1), outline ideas for stronger attacks against KAD (Section 6.2), and briefly mention directions for developing a more robust detection mechanism (Section 6.3).
6.1 Limitations of improving KAD
@@
1. Adversarial training of detection LLM $\boldsymbol{g}$. As discussed in Section 3.4 and noted by prior work [16], adversarial training can help $g$ resist specific adaptive attacks that trigger Type II failures (described in Section 3.4). However, there exists a practically unbounded set of semantically equivalent injected prompts that can induce the detection LLM to follow the detection instruction; many of these injected prompts may not be encountered during training. Generalizing across this space is inherently hard. While full generalization is a challenge for any defense, KAD is particularly brittle because its failure arises from a structural flaw: it expects the detection LLM to follow injected instructions (Axiom 2 in Section 3.3), enabling systematic adversarial attacks without optimization.
2. Obscuring the detection instruction. Our DataFlip attack leverages the knowledge of the detection instruction’s structure, but hiding the detection instruction does not eliminate the problem. An adaptive adversary can still craft contaminated inputs that induce the detection LLM to behave differently depending on whether it receives the detection or target instruction—following the detection instruction in one case and the injected instruction in the other. This is always possible because the detection and target instruction are inherently distinct, one expects a predetermined answer and the other depends on the input. As LLMs improve at instruction-following, such behavioral bifurcation becomes increasingly achievable, even without access to the detection instruction.
@@
6.2 Stronger attack against KAD
Our proposed DataFlip attack is handcrafted, leveraging simple IF/ELSE semantics to induce adversarial behavior. This could be further strengthened by incorporating optimization-based triggers—such as those used in Universal [25] or NeuralExec [24] attacks—that search for distinct trigger patterns. These triggers can be tailored to selectively increase the likelihood that the detection LLM follows the IF branch while the backend LLM follows the ELSE branch, thereby improving attack efficiency.
6.3 Towards ideal detection mechanism
@@
7 Conclusion
Acknowledgments
Sarthak Choudhary, Divyam Anshumaan, Nils Palumbo, and Somesh Jha are partially supported by DARPA under agreement number 885000, NSF CCF-FMiTF-1836978 and ONR N00014-21-1-2492.

### v3

@@ 5.4. RQ3: KAD vs Instruction Fine-tuned FNR
Design. We quantitatively measure the advantage (if any) of the fine-tuning process used in Strong KAD defenses such as DataSentinel by comparing the false negative rate (FNR) of the detection model $g$ with an instruction fine-tuned version of Mistral-7B [^3] https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, under two adversarial settings: Combined Attack and DataFlip.
@@
Setup. We consider contaminated inputs crafted using Combined Attack and DataFlip on the Prompt Injection dataset from Section 5.1.2. We report FPRs and FNRs (Section 5.1.3) corresponding to each attack, referred to as FNRCombined and FNRDataFlip, for the DataSentinel fine-tuned detection LLM and Mistral-Instruct.
Results. We see in Table 2 that, despite fine-tuning on samples crafted using the adaptive attacks, the FNRCombined of the DataSentinel detector is almost identical to the Mistral-Instruct model. Furthermore, we find that FNRDataFlip of the DataSentinel detector is slightly worse than Mistral-Instruct in most categories, with Grammar Correction in particular having a $83\%$ gap in performance.
@@
In fact, we note that the values of FNRCombined and FNRDataFlip both indicate a high tendency to follow the injected instruction as per our analysis in Section 3. The Combined Attack takes no advantage of the fact that the detection LLM is attempting to follow the instruction, resulting in the secret key not being produced. However, when we switch to DataFlip, instruction following results in the extraction of the secret key, yielding high FNR. This once again underscores the flawed nature of the KAD objective and the negative impact of Strong KAD defenses which optimize for it.
6. Discussions
We discuss the limitations of potential improvements to KAD, stronger attack designs, and directions for robust detection in Appendix A. In brief, modest enhancements such as adversarial training (liu2025datasentinel), obscuring detection instructions, or isolating injected content fail to address KAD’s structural vulnerability (Axiom 2 in Section 3.3), leaving it open to systematic exploitation (Section 3.4). We also outline how optimization-based triggers from prior work (liu2024automatic; pasquini2024neural) could strengthen attacks, and argue that interpretability-driven approaches (e.g., attention analysis, influence tracing) offer a more principled path to robust detection.
7. Conclusion
Acknowledgements.
We thank the anonymous reviewers for their valuable feedback and Neil Zhenqiang Gong along with the other authors of DataSentinel for providing detailed clarifications of their work, which enabled a consistent evaluation of our attack. Sarthak Choudhary, Divyam Anshumaan, Nils Palumbo, and Somesh Jha are partially supported by DARPA under agreement number 885000, NSF CCF-FMiTF-1836978 and ONR N00014-21-1-2492.
Appendix A Additional Discussions
@@
In this section, we discuss the limitations of potential attempts to improve KAD (Section A.1), outline ideas for stronger attacks against KAD (Section A.2), and briefly mention directions for developing a more robust detection mechanism (Section A.3).
A.1. Limitations of improving KAD
@@
1. Adversarial training of detection LLM $\boldsymbol{g}$. As discussed in Section 3.4 and noted by prior work (liu2025datasentinel), adversarial training can help $g$ resist specific adaptive attacks that trigger Type II failures (described in Section 3.4). However, there exists a practically unbounded set of semantically equivalent injected prompts that can induce the detection LLM to follow the detection instruction; many of these injected prompts may not be encountered during training. Generalizing across this space is inherently hard. While full generalization is a challenge for any defense, KAD is particularly brittle because its failure arises from a structural flaw: it expects the detection LLM to follow injected instructions (Axiom 2 in Section 3.3), enabling systematic adversarial attacks without optimization.
2. Obscuring the detection instruction. Our DataFlip attack leverages the knowledge of the detection instruction’s structure, but hiding the detection instruction does not eliminate the problem. An adaptive adversary can still craft contaminated inputs that induce the detection LLM to behave differently depending on whether it receives the detection or target instruction—following the detection instruction in one case and the injected instruction in the other. This is always possible because the detection and target instruction are inherently distinct; one expects a predetermined answer, and the other depends on the input. As LLMs improve at instruction-following, such behavioral bifurcation becomes increasingly achievable, even without access to the detection instruction.
@@
A.2. Stronger attack against KAD
Our proposed DataFlip attack is handcrafted, leveraging simple IF/ELSE semantics to induce adversarial behavior. This could be further strengthened by incorporating optimization-based triggers—such as those used in Universal (liu2024automatic) or NeuralExec (pasquini2024neural) attacks—that search for distinct trigger patterns. These triggers can be tailored to selectively increase the likelihood that the detection LLM follows the IF branch while the backend LLM follows the ELSE branch, thereby improving attack efficiency.
A.3. Towards ideal detection mechanism